Gary Company purchases sugar beets for $60 a batch. The total cost to crush a batch of sugar beets is $12. At that point there are two intermediate products - beet fiber and beet juice. The beet fiber can be sold as is for $28 or processed further at a cost of $20 to make industrial fiber that can be sold for $67. The beet juice can be sold as is for $39 or processed further for $24 to make refined sugar that is sold for $54. Which of the products - beet fiber and/or beet juice - should be processed further? Explain why including computations.
Answer #1

Beet Fiber

Selling price of beet fiber at point of separation is 28.

Cost for further processing 20 and selling price will be 67.

Net benefit = Selling price - Further processing cost - Opportunity cost

   = 67-20-28 = 19

It is not advisable to further process to Beet Fiber.

Beet Juice

Selling price of Beet Juice at point of separation is 39.

Cost for further processing 24 and selling price of refined sugar will be 54.

Net benefit = Selling price - Further processing cost - Opportunity cost

   = 54-24-39 = (9) loss

It is not advisable to further process to Beet Juice also.